OpenStudy (anonymous):

Square 6x^4y^3 - square 24x^6y Square 6x^4y^3 - square 24x^6y @Mathematics

OpenStudy (anonymous):

\[\sqrt{6x^4y^3}-\sqrt{24x^6y}\]

OpenStudy (goldrush18):

=(6x^4y^3)^1/2 - (24x^6y)^1/2 =6^1*1/2 x^4*1/2 y^3*1/2 - 24^1*1/2 x^6*1/2 y^1*1/2 =6^1/2 x^2 y ^3/2 - 24^1/2 x^3 y^1/2 =3 x^2 y^3/2 - 12 x^3 y^1/2

OpenStudy (goldrush18):

hope u understand so far im not sure yet how to finish it whether i should divide im still thinking

OpenStudy (anonymous):

I'm a bit confused :/

OpenStudy (goldrush18):

where are u confused